Given the linear equation y=-7/8x+2, find the y-coordinates of the points (-8,), (0,), and (8,). Please show all of your work
Answer by mananth(16949)

Plug the x values in the equation y=mx+b , to generate y values y= 0.88 x + 2

x= -8 , y= 0.88 * -8 + 2 , y= -7 + 2 , y= -5 ,
x= 0 , y= 0.88 * 0 + 2 , y= 0 + 2 , y= 2
x= 8 , y= 0.88 * 8 + 2 , y= 7 + 2 , y= 9